# Unbiased Regression-Adjusted Estimation of Average Treatment Effects in Randomized Controlled Trials

### v1

@@ 7 Conclusion
This paper develops leave-one-out regression-adjusted estimators for treatment effects under simple and complete random assignment. We show that these estimators are exactly unbiased in finite populations, achieve the asymptotic efficiency of the interacted regression of lin2013agnostic, and admit closed-form expressions for their variance. Regularization through leave-one-out ridge adjustment ensures robustness to high-leverage observations and stabilizes inference even in small samples. Our theoretical results establish a design-based foundation for regression adjustment that unifies classical unbiasedness, modern efficiency results, and practical inference procedures. Empirical evaluations confirm the estimators’ excellent finite-sample performance across different designs.
Beyond the settings considered here, future work could extend our framework to stratified or covariate-adaptive designs, to clustered or networked experiments, and to high-dimensional or nonparametric adjustments. Such extensions would further connect design-based causal inference with modern statistical learning while preserving the transparency and interpretability that motivate experimental design in the first place.
Dataset Method Bias STD RMSE CI coverage (95%) CI average length (95%)
Statehood REG 0.0432 0.5100 0.5118 0.7259 0.5785
Ridge-REG 0.0559 0.2692 0.2749 0.8989 0.4515
LOORA-HT 0.0009 0.2856 0.2856 0.9531 0.5901
LOORA-DM 0.0017 0.2781 0.2781 0.9557 0.5849
Lightbulb REG 0.0023 0.1048 0.1048 0.9094 0.1783
Ridge-Reg 0.0297 0.0861 0.0911 0.9190 0.1533
LOORA-HT 0.0000 0.0822 0.0822 0.9616 0.1705
LOORA-DM 0.0000 0.0797 0.0797 0.9623 0.1671
Table 3: Performance metrics for four methods on two datasets with treatment assignment probabilities all equal to $0.5$.
@@ Appendix A Linear Algebra Tools
Lemma A.1 (alaoui2015fast; fahrbach2022subquadratic).
Proof of Lemma˜2.
Let ${\bm{U}}{\boldsymbol{\Sigma}}{\bm{V}}$ be the compact SVD of $\bm{X}$} and $r$ be its rank. Then by Lemma˜A.1,
Since $\bm{U}$} is an orthogonal matrix, for all $i\in[n]$,
$\sum_{j=1}^{r}u_{ij}^{2}\leq 1.$
Without loss of generality suppose $\sum_{j=1}^{r}u_{ij}^{2}=1$ since scaling up $u_{ij}$’s only increases the right-hand side of (21). By definition of $\zeta$, for all $i\in[n]$,
Therefore by (21),
Let $Y$ be a random variable that is equal to $\sigma_{j}^{2}$ with probability $u_{ij}^{2}$ and $\phi:{\mathbb{R}}_{\geq 0}\to{\mathbb{R}}_{\geq 0}$ be a function with $\phi(g)=\frac{g}{g+c\ell_{i}}$. Therefore
Lemma A.2 ((Lemma 3.2 of miller1974unbalanced)).
@@
Proof of Lemma˜B.1.
Let
[…]
One can observe that for any $\bm{b}$},
[…]
Therefore it immediately follows by Lemma˜A.2 that
${\bm{x}}_{i}^{\top}\boldsymbol{\beta}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}^{(-i)}=\frac{{\bm{x}}_{i}^{\top}({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{x}}_{i}(y_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta})}{1-h_{i}({\bm{X}},\lambda)}=\frac{h_{i}({\bm{X}},\lambda)(y_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta})}{1-h_{i}({\bm{X}},\lambda)}$
$\implies$ ${\bm{x}}_{i}^{\top}\boldsymbol{\beta}-y_{i}=(1-h_{i}({\bm{X}},\lambda))({\bm{x}}_{i}^{\top}\boldsymbol{\beta}^{(-i)}-y_{i})\,.$
Lemma A.3.
Let $\lambda^{(1)}\leq\lambda^{(2)}$ and for $j\in\{1,2\}$, ${\bm{v}}\in{\mathbb{R}}^{n}$, and
$\boldsymbol{\beta}_{\lambda^{(j)}}=\mathop{\mathrm{argmin}\,}\limits_{\boldsymbol{\beta}}\left\|{\bm{X}}\boldsymbol{\beta}-{\bm{v}}\right\|_{2}^{2}+\lambda^{(j)}\left\|\boldsymbol{\beta}\right\|_{2}^{2}.$
Then $\left\|{\bm{X}}\boldsymbol{\beta}_{\lambda^{(1)}}-{\bm{v}}\right\|_{2}^{2}\leq\left\|{\bm{X}}\boldsymbol{\beta}_{\lambda^{(2)}}-{\bm{v}}\right\|_{2}^{2}.$
@@
Let ${\bm{U}}{\boldsymbol{\Sigma}}{\bm{V}}^{\top}$ be the SVD of $\bm{X}$}. Then
${\bm{X}}({\bm{X}}^{\top}{\bm{X}}+\lambda^{(j)}{\bm{I}})^{-1}{\bm{X}}^{\top}={\bm{U}}{\boldsymbol{\Sigma}}({\boldsymbol{\Sigma}}^{\top}{\boldsymbol{\Sigma}}+\lambda^{(j)}{\bm{I}})^{-1}{\boldsymbol{\Sigma}}^{\top}{\bm{U}}^{\top}$
Therefore
$({\bm{I}}-{\bm{X}}({\bm{X}}^{\top}{\bm{X}}+\lambda^{(j)}{\bm{I}})^{-1}{\bm{X}}^{\top})^{2}={\bm{U}}({\bm{I}}-{\boldsymbol{\Sigma}}({\boldsymbol{\Sigma}}^{\top}{\boldsymbol{\Sigma}}+\lambda^{(j)}{\bm{I}})^{-1}{\boldsymbol{\Sigma}}^{\top})^{2}{\bm{U}}^{\top}$
Note ${\boldsymbol{\Sigma}}^{\top}{\boldsymbol{\Sigma}}+\lambda^{(1)}{\bm{I}}\preceq{\boldsymbol{\Sigma}}^{\top}{\boldsymbol{\Sigma}}+\lambda^{(2)}{\bm{I}}$. Therefore
${\bm{I}}-{\boldsymbol{\Sigma}}({\boldsymbol{\Sigma}}^{\top}{\boldsymbol{\Sigma}}+\lambda^{(1)}{\bm{I}})^{-1}{\boldsymbol{\Sigma}}^{\top}\preceq{\bm{I}}-{\boldsymbol{\Sigma}}({\boldsymbol{\Sigma}}^{\top}{\boldsymbol{\Sigma}}+\lambda^{(2)}{\bm{I}})^{-1}{\boldsymbol{\Sigma}}^{\top}.$
Now since both of these matrices are diagonal, we have
$({\bm{I}}-{\boldsymbol{\Sigma}}({\boldsymbol{\Sigma}}^{\top}{\boldsymbol{\Sigma}}+\lambda^{(1)}{\bm{I}})^{-1}{\boldsymbol{\Sigma}}^{\top})^{2}\preceq({\bm{I}}-{\boldsymbol{\Sigma}}({\boldsymbol{\Sigma}}^{\top}{\boldsymbol{\Sigma}}+\lambda^{(2)}{\bm{I}})^{-1}{\boldsymbol{\Sigma}}^{\top})^{2}.$
The result then follows by observing that
$\left\|{\bm{X}}\boldsymbol{\beta}^{(1)}-{\bm{y}}\right\|_{2}^{2}={\bm{y}}^{\top}({\bm{I}}-{\bm{X}}({\bm{X}}^{\top}{\bm{X}}+\lambda^{(j)}{\bm{I}})^{-1}{\bm{X}}^{\top})^{2}{\bm{y}}$
@@ Appendix B Main Proofs
LOORA-HT depends on a set of $n$ leave-one-out regression coefficients $\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}$. Next lemma characterizes the difference between the full-sample regularized least-squares coefficient vector and its leave-one-out counterparts (see miller1974unbalanced for a simpler version for least-squares).
Lemma B.1.
@@
Lemma˜B.1 basically gives us the intuition on the deviation that will happen by solving the left-one-out regression for the quantities such coefficients of the regression, value of errors, and estimations for each of the entries $i\in[n]$. The key point is that all the deviations depend on the value of the leverage score of the removed row, which further helps us characterize the robustness of our leave-one-out estimators to observation removal in inference time, which is a point of concern, especially in small sample settings where a single observation could have a large (close to $1$) leverage score that makes them affect quantities such as variance and confidence intervals to a large extent, which is not desirable. Hence, this is one of the first places that can motivate us to favor regularized regression adjustment methods, since as we will see in Lemma˜2, it will help us be able to uniformly decrease the leverage scores across all observations.
Lemma B.2.
@@
where $\boldsymbol{\beta}_{\lambda}$ is defined as in Equation˜6. Therefore,
@@ B.1 Omitted Proofs of Section˜3
As a result, $\widehat{\tau}_{\textup{LHT}}$ is unbiased. By Lemma˜B.2
@@
By Lemma˜B.2, we have $\widehat{\tau}-\tau=\frac{1}{n}{\bm{z}}^{\top}\boldsymbol{g}$. Therefore
@@
${\mathbb{E}}\Bigg[\frac{z_{i}(r_{i}\mu_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})}{q_{i}(1-\widetilde{h}_{\lambda ii})}$ $\frac{z_{j}{\bm{x}}_{i}^{\top}(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}_{-j}^{\top}({\bm{z}}{\bm{t}}/\boldsymbol{q})_{-j}}{q_{j}(1-\widetilde{h}_{\lambda jj})}\Bigg]$
$={\mathbb{E}}\Bigg[\frac{z_{j}}{q_{j}}\Bigg]{\mathbb{E}}\Bigg[\frac{z_{i}(r_{i}\mu_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})}{q_{i}(1-\widetilde{h}_{\lambda ii})}\frac{{\bm{x}}_{i}^{\top}(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}_{-j}^{\top}({\bm{z}}{\bm{t}}/\boldsymbol{q})_{-j}}{(1-\widetilde{h}_{\lambda jj})}\Bigg]=0.$
@@
${\mathbb{E}}\left[\frac{z_{i}(r_{i}\mu_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta})}{q_{i}(1-\widetilde{h}_{\lambda ii})}\frac{z_{j}(r_{j}\mu_{j}-{\bm{x}}_{j}^{\top}\boldsymbol{\beta})}{q_{j}(1-\widetilde{h}_{\lambda jj})}\right]=0.$
@@
AA about here.
Proof of Lemma˜3.
First, observe that
@@
$\boldsymbol{\beta}^{*}_{\lambda}=\lim_{n\to\infty}(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}^{\top}{\boldsymbol{\mu}}.$
Under Assumption˜2, $(n^{-1}{\bm{X}}^{\top}{\bm{X}})\to\Sigma_{X}$, where $\Sigma_{X}$ is bounded and positive definite. Moreover, by Assumption˜3, since $m=\inf_{i\in\mathbb{N}}\min\{p_{i},1-p_{i}\}>0$, we have
$n^{-1}\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}=n^{-1}{\bm{X}}^{\top}{\bm{R}}^{-2}{\bm{X}}\to\Sigma_{\widetilde{X}}$,
where $\Sigma_{\widetilde{X}}$ is bounded and invertible. Hence,
$\left\|(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\right\|_{2}=O(n^{-1}).$ (22)
Furthermore,
$\widetilde{{\bm{X}}}^{\top}{\boldsymbol{\mu}}=\sum_{i=1}^{n}{\bm{x}}_{i}\left(\frac{y_{i}^{(1)}}{p_{i}}+\frac{y_{i}^{(0)}}{1-p_{i}}\right),$
and by Assumptions˜2 and 3, $\left\|\widetilde{{\bm{X}}}^{\top}{\boldsymbol{\mu}}\right\|_{2}=O(n)$. It follows that the limit $(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}^{\top}{\boldsymbol{\mu}}$ exists and that $\boldsymbol{\beta}^{*}_{\lambda}$ is finite.
Step 1: Consistency of the full-sample estimator. Define
$\widehat{\boldsymbol{\beta}}_{\lambda}=(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{y}}}$.
By construction, $\widetilde{{\bm{y}}}={\boldsymbol{\mu}}+{\bm{z}}{\bm{t}}/\boldsymbol{q}$, where ${\mathbb{E}}[z_{i}t_{i}/q_{i}]=0$ for all $i$. Because ${\bm{x}}_{i}$, $t_{i}$, and $1/p_{i},1/(1-p_{i})$ are uniformly bounded, the weak law of large numbers implies
$\frac{1}{n}\widetilde{{\bm{X}}}^{\top}({\bm{z}}{\bm{t}}/\boldsymbol{q})\xrightarrow{p}\boldsymbol{0}.$
By (22) and Slutsky’s theorem,
$(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{y}}}\xrightarrow{p}(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}^{\top}{\boldsymbol{\mu}},$
and hence $\widehat{\boldsymbol{\beta}}_{\lambda}\xrightarrow{p}\boldsymbol{\beta}^{*}_{\lambda}$.
Step 2: Asymptotic equivalence of leave-one-out and full-sample estimators. By Lemma˜B.1,
@@
$\widehat{\boldsymbol{\beta}}_{\lambda}-\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}=\frac{(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{x}}}_{i}(\widetilde{y}_{i}-\widetilde{{\bm{x}}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda})}{1-\widetilde{h}_{\lambda ii}},$ (23)
where $\widetilde{h}_{\lambda ii}=\widetilde{{\bm{x}}}_{i}^{\top}(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{x}}}_{i}$. From Assumptions˜1 and 3 and (22),
$\widetilde{h}_{\lambda ii}\leq\|\widetilde{{\bm{x}}}_{i}\|_{2}^{2}\,\left\|(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\right\|_{2}=O(n^{-1}),$
so $\widetilde{h}_{\lambda ii}\to 0$ uniformly in $i$, and the denominator in (23) converges to $1$.
Next, by Assumption˜1 and the inequalities $\|{\bm{y}}^{(1)}\|_{\infty}\leq(nL)^{1/4}$ and $\|{\bm{y}}^{(0)}\|_{\infty}\leq(nL)^{1/4}$, the responses are uniformly bounded in $O(n^{1/4})$. Because $\|(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\|_{2}=O(n^{-1})$ and $\|\widetilde{{\bm{x}}}_{i}\|_{2}=O(1)$, it follows from (23) that
@@
$\left\|\widehat{\boldsymbol{\beta}}_{\lambda}-\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}\right\|_{2}=O(n^{-3/4}).$
Hence $\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}\xrightarrow{p}\widehat{\boldsymbol{\beta}}_{\lambda}$ for each $i$. Combining Steps 1 and 2,

### v2

@@ 7 Conclusion
This article develops leave-one-out regression-adjusted estimators of average treatment effects under simple and complete random assignment. We show that these estimators are exactly unbiased in finite populations, achieve the asymptotic efficiency of the interacted regression of Lin (2013), and admit closed-form expressions for their variance. Regularization through leave-one-out ridge adjustment ensures robustness to high-leverage observations and stabilizes inference even in small samples. Our theoretical results establish a design-based foundation for regression adjustment that unifies classical unbiasedness, modern efficiency results, and practical inference procedures. Empirical evaluations confirm the estimators’ excellent finite-sample performance across different designs.
Natural next steps include extending the framework to stratified or covariate-adaptive designs and to clustered or networked experiments, and incorporating high-dimensional or nonparametric adjustments. Taken together, these directions define a broad agenda for design-based regression adjustment in modern experimental settings.
@@ Appendix A Linear Algebra Tools
Lemma 3 ((Lemma 3.2 of Miller (1974))).
@@
Lemma 4 (Alaoui and Mahoney (2015); Fahrbach et al. (2022)).
Proof of Lemma˜1.
Let ${\bm{U}}{\boldsymbol{\Sigma}}{\bm{V}}$ be the compact SVD of $\bm{X}$} and $r$ be its rank. Then by Lemma˜4,
Since $\bm{U}$} is an orthogonal matrix, for all $i\in[n]$,
$\sum_{j=1}^{r}u_{ij}^{2}\leq 1.$
By definition of $\zeta$, for all $i\in[n]$,
Therefore by (14),
Let $Y$ be a random variable that is equal to $\sigma_{j}^{2}$ with probability $u_{ij}^{2}$, and take value 0 with probability ($1-\sum_{j=1}^{r}u^{2}_{ij}$), and $\phi:{\mathbb{R}}_{\geq 0}\to{\mathbb{R}}_{\geq 0}$ be a function with $\phi(g)=\frac{g}{g+c\ell_{i}}$. Therefore
Lemma 5.
Let $\boldsymbol{M}\in{\mathbb{R}}^{n\times k}$ and ${\boldsymbol{\mu}}\in{\mathbb{R}}^{n}$. Then
$\min_{{\bm{b}}\in{\mathbb{R}}^{k}}\left\|(\boldsymbol{M}-\overline{\boldsymbol{M}}){\bm{b}}-({\boldsymbol{\mu}}-\overline{{\boldsymbol{\mu}}})\right\|_{2}^{2}=\min_{{\bm{b}}\in{\mathbb{R}}^{k+1}}\left\|\begin{bmatrix}\boldsymbol{M}&~~\boldsymbol{1}\end{bmatrix}{\bm{b}}-{\boldsymbol{\mu}}\right\|_{2}^{2}.$
@@
Let $\boldsymbol{N}=\begin{bmatrix}(\boldsymbol{M}-\overline{\boldsymbol{M}})&~~\boldsymbol{1}\end{bmatrix}$ and let $\boldsymbol{H}^{(\boldsymbol{N})}$ denote the projection matrix onto the column space of $\boldsymbol{N}$, i.e., $\boldsymbol{H}^{(\boldsymbol{N})}=\boldsymbol{N}(\boldsymbol{N}^{\top}\boldsymbol{N})^{-1}\boldsymbol{N}^{\top}$. Then
$\min_{{\bm{b}}\in{\mathbb{R}}^{k+1}}\left\|\boldsymbol{N}{\bm{b}}-{\boldsymbol{\mu}}\right\|_{2}^{2}=\left\|\boldsymbol{H}^{(\boldsymbol{N})}{\boldsymbol{\mu}}-{\boldsymbol{\mu}}\right\|_{2}^{2}.$
Since $\boldsymbol{1}$ is orthogonal to the columns of $\boldsymbol{M}-\overline{\boldsymbol{M}}$, we have
$\boldsymbol{H}^{(\boldsymbol{N})}{\boldsymbol{\mu}}=\boldsymbol{H}^{(\boldsymbol{M}-\overline{\boldsymbol{M}})}{\boldsymbol{\mu}}+\boldsymbol{H}^{(\boldsymbol{1})}{\boldsymbol{\mu}}=\boldsymbol{H}^{(\boldsymbol{M}-\overline{\boldsymbol{M}})}({\boldsymbol{\mu}}-\overline{{\boldsymbol{\mu}}})+\overline{{\boldsymbol{\mu}}},$ (15)
where the second equality uses $\overline{{\boldsymbol{\mu}}}=\boldsymbol{H}^{(\boldsymbol{1})}{\boldsymbol{\mu}}$ and $\boldsymbol{H}^{(\boldsymbol{M}-\overline{\boldsymbol{M}})}\overline{{\boldsymbol{\mu}}}=0$. Therefore,
$\left\|\boldsymbol{H}^{(\boldsymbol{N})}{\boldsymbol{\mu}}-{\boldsymbol{\mu}}\right\|_{2}^{2}$ $=\left\|\boldsymbol{H}^{(\boldsymbol{M}-\overline{\boldsymbol{M}})}({\boldsymbol{\mu}}-\overline{{\boldsymbol{\mu}}})-({\boldsymbol{\mu}}-\overline{{\boldsymbol{\mu}}})\right\|_{2}^{2}$
$=\min_{{\bm{b}}\in{\mathbb{R}}^{k}}\left\|(\boldsymbol{M}-\overline{\boldsymbol{M}}){\bm{b}}-({\boldsymbol{\mu}}-\overline{{\boldsymbol{\mu}}})\right\|_{2}^{2}.$
Finally note that the span of columns of $\begin{bmatrix}\boldsymbol{M}&~~\boldsymbol{1}\end{bmatrix}$ and $\boldsymbol{N}$ are the same. Thus
$\min_{{\bm{b}}\in{\mathbb{R}}^{k+1}}\left\|\boldsymbol{N}{\bm{b}}-{\boldsymbol{\mu}}\right\|_{2}^{2}=\min_{{\bm{b}}\in{\mathbb{R}}^{k+1}}\left\|\begin{bmatrix}\boldsymbol{M}&~~\boldsymbol{1}\end{bmatrix}{\bm{b}}-{\boldsymbol{\mu}}\right\|_{2}^{2},$
which concludes the result. ∎
@@ Appendix B Main Proofs
LOORA-HT depends on a set of $n$ leave-one-out regression coefficients $\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}$. Next lemma characterizes the difference between the full-sample regularized least-squares coefficient vector and its leave-one-out counterparts (see Lemma˜3 for a simpler version for least-squares).
Lemma 6.
@@
Proof.
Let
[…]
One can observe that for any $\bm{b}$},
[…]
Therefore it immediately follows by Lemma˜3 that
$\widehat{\boldsymbol{\beta}}_{\lambda}-\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}=\frac{({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{x}}_{i}(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda})}{1-h_{\lambda ii}},$ (16)
and therefore
${\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}=\frac{{\bm{x}}_{i}^{\top}({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{x}}_{i}(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda})}{1-h_{i}({\bm{X}},\lambda)}=\frac{h_{i}({\bm{X}},\lambda)(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda})}{1-h_{i}({\bm{X}},\lambda)}$
$\implies$ ${\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}-y_{i}=(1-h_{i}({\bm{X}},\lambda))({\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}-y_{i})\,.$
Thus,
${\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}$ $={\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}-\frac{h_{\lambda ii}(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda})}{1-h_{\lambda ii}}=\frac{{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}-h_{\lambda ii}y_{i}}{1-h_{\lambda ii}}$
$=\frac{{\bm{x}}_{i}^{\top}({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{X}}^{\top}{\bm{y}}-{\bm{x}}_{i}^{\top}({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{x}}_{i}y_{i}}{1-h_{\lambda ii}}$
$=\frac{{\bm{x}}_{i}^{\top}({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{X}}_{-i}^{\top}{\bm{y}}_{-i}}{1-h_{\lambda ii}}.$
Lemma˜6 characterizes the deviation of the solution, estimator, and residual when we solve a leave-one-out regression instead of the full regression. The key observation is that all deviations depend on the leverage score of the removed row, which in turn helps characterize the robustness of our leave-one-out estimators to observation removal at inference time. This issue is particularly important in small-sample settings, where a single observation may have a large (close to $1$) leverage score and thus exert a disproportionate influence on quantities such as variance estimates and confidence intervals, which is undesirable. Hence, this result provides one of the first motivations for favoring regularized regression adjustment methods: as seen in Lemma˜1, regularization enables us to uniformly reduce leverage scores across all observations.
Lemma 7.
@@
where $\boldsymbol{\beta}_{\lambda}$ is defined as in Equation˜4. Therefore,
@@ B.1 Omitted Proofs of Section˜3
As a result, $\widehat{\tau}_{\textup{LHT}}$ is unbiased. By Lemma˜7
@@
By Lemma˜7, we have $\widehat{\tau}-\tau=\frac{1}{n}{\bm{z}}^{\top}\boldsymbol{g}$. Therefore
@@
${\mathbb{E}}\Bigg[\frac{z_{i}(r_{i}\mu_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})}{q_{i}(1-\widetilde{h}_{\lambda ii})}$ $\frac{z_{j}{\bm{x}}_{j}^{\top}(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}_{-j}^{\top}({\bm{z}}{\bm{t}}/\boldsymbol{q})_{-j}}{q_{j}(1-\widetilde{h}_{\lambda jj})}\Bigg]$
$={\mathbb{E}}\Bigg[\frac{z_{j}}{q_{j}}\Bigg]{\mathbb{E}}\Bigg[\frac{z_{i}(r_{i}\mu_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})}{q_{i}(1-\widetilde{h}_{\lambda ii})}\frac{{\bm{x}}_{j}^{\top}(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}_{-j}^{\top}({\bm{z}}{\bm{t}}/\boldsymbol{q})_{-j}}{(1-\widetilde{h}_{\lambda jj})}\Bigg]=0.$
@@
${\mathbb{E}}\left[\frac{z_{i}(r_{i}\mu_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})}{q_{i}(1-\widetilde{h}_{\lambda ii})}\frac{z_{j}(r_{j}\mu_{j}-{\bm{x}}_{j}^{\top}\boldsymbol{\beta}_{\lambda})}{q_{j}(1-\widetilde{h}_{\lambda jj})}\right]=0.$
@@
Proof of Theorem˜2.
@@
We first consider the case where $\left\|\widetilde{{\bm{X}}}\boldsymbol{\beta}_{\lambda}-{\boldsymbol{\mu}}\right\|_{2}\to\infty$. We write
@@
$\widehat{\tau}_{\textup{LHT}}^{*}$ $=\frac{1}{n}\sum_{i=1}^{n}\frac{z_{i}}{q_{i}}\bigl(y_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}\bigr),~~\widehat{\tau}_{\textup{LHT}}=\frac{1}{n}\sum_{i=1}^{n}\frac{z_{i}}{q_{i}}\bigl(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}\bigr)$,
and decompose
@@
$\frac{\sqrt{n}(\widehat{\tau}_{\textup{LHT}}-\tau)}{\left\|\widetilde{{\bm{X}}}\boldsymbol{\beta}_{\lambda}-{\boldsymbol{\mu}}\right\|_{2}/\sqrt{n}}=A_{n}+B_{n},$
where